# Probabilistic Time Integration for Semi-explicit PDAEs

### v1

@@ 4.2. Probabilistic midpoint scheme
The (deterministic) midpoint scheme applied to the PDAE (2.1) results in
@@
(4.3b) $\mathcal{B}u^{n+1}\hskip 80.52135pt$ $=\phantom{\tau}g^{n+1}\hskip 24.75375pt\text{in }\mathcal{Q}^{*},$
where $f^{n+\frac{1}{2}}=f\left(t^{n}+\frac{\tau}{2},\frac{1}{2}(u^{n}+u^{n+1})\right)$. The result is a second-order scheme, i.e., we get (2.7) with $q=2$. Using again the fact that $u^{n+1}_{c}=\mathcal{B}^{-}g^{n+1}$ and restricting equation (4.3a) to test functions in $\mathcal{V}_{\ker}$, we obtain
$\big{(}\operatorname{id}+\tfrac{\tau}{2}\mathcal{A}\big{)}\,u_{\ker}^{n+1}=\big{(}\operatorname{id}-\tfrac{\tau}{2}\mathcal{A}\big{)}\,u^{n}_{\ker}+\tau f\big{(}t^{n}+\tfrac{\tau}{2},\tfrac{1}{2}(u^{n}+u^{n+1})\big{)}-\mathcal{B}^{-}\big{(}g^{n+1}-g^{n}\big{)}\quad\text{in }\mathcal{V}_{\ker}^{*}$
@@
$\Psi_{\tau}(t,v)=\big{(}\operatorname{id}-\tfrac{\tau}{2}\mathcal{A}\big{)}\,v-\tfrac{\tau}{2}\mathcal{A}\Psi_{\tau}(t,v)+\tau f\big{(}t+\tfrac{\tau}{2},v^{*}(t)\big{)}-\mathcal{B}^{-}\big{(}g(t+\tau)-g(t)\big{)}$
with $v^{*}(t)\coloneqq\frac{1}{2}\big{(}v+\mathcal{B}^{-}g(t)+\Psi_{\tau}(t,v)+\mathcal{B}^{-}g(t+\tau)\big{)}$.
@@
Given $v_{1},v_{2}\in\mathcal{V}_{\ker}$, we define $w_{0}$ and $w_{1}$ as in the proof of Lemma 4.1. It directly follows that $v_{1}^{*}-v_{2}^{*}=\frac{1}{2}(w_{0}+w_{1})$ and, in turn, leads to
@@
$\qquad\qquad+\Big{(}f\big{(}t+\tfrac{\tau}{2},v_{1}^{*}\big{)}-f\big{(}t+\tfrac{\tau}{2},v_{2}^{*}\big{)},w_{0}+w_{1}\Big{)}_{\mathcal{H}}$
@@
$\begin{bmatrix}\operatorname{id}+\frac{\tau}{2}\mathcal{A}&\tau\mathcal{B}^{*}\\ \mathcal{B}&\end{bmatrix}\begin{bmatrix}U^{n+1}\\ \lambda^{n+1}\end{bmatrix}=\begin{bmatrix}(\operatorname{id}-\frac{\tau}{2}\mathcal{A})U^{n}+\tau f^{n+\frac{1}{2}}+\xi^{n}(\tau)\\ g^{n+1}\end{bmatrix},$
@@ 4.3. Probabilistic exponential Euler scheme
As third example, we derive the probabilistic version of the exponential Euler method. The corresponding deterministic scheme for constrained systems has been introduced and examined in [AZ20]. Recall that $\mathcal{A}_{\ker}$ denotes the restriction of $\mathcal{A}$ to the kernel of the constraint operator $\mathcal{B}$. Moreover, we introduce $\mathcal{H}_{\ker}$ as the closure of $\mathcal{V}_{\ker}$ in $\mathcal{H}$. Due to Assumption 2.2, we know that $-\mathcal{A}_{\ker}$ generates an analytic semigroup on $\mathcal{H}_{\ker}$; see [Paz83, Ch. 7, Th. 2.7].
@@
$u_{\ker}(t^{n+1})$ $=u(t^{n+1})-\mathcal{B}^{-}g^{n+1}$
(4.4) $=e^{-\tau\mathcal{A}_{\ker}}\big{[}u(t^{n})-\mathcal{B}^{-}g^{n}\big{]}+\int_{t^{n}}^{t^{n+1}}e^{-(t^{n+1}-s)\mathcal{A}_{\ker}}\iota_{0}\big{[}f(s,u(s))-\mathcal{B}^{-}\dot{g}(s)\big{]}\,\mathrm{d}s$
with $\iota_{0}\colon\mathcal{H}\equiv\mathcal{H}^{*}\rightarrow\mathcal{H}^{*}_{\ker}\equiv\mathcal{H}_{\ker}$ denoting the restriction of test functions (which will be omitted from now on). Yet another important tool for the construction are the recursively defined $\varphi$-functions,
@@
with $\varphi_{k}(0)=1/k!$ for $z=0$. Since $-\mathcal{A}_{\ker}$ generates an analytic semigroup, it follows from [HO10, Lem. 2.4] that $\varphi_{k}(-\tau\mathcal{A}_{\ker})\colon\mathcal{H}_{\ker}\to\mathcal{H}_{\ker}$ defines a bounded operator.
@@
(4.6) $u^{n+1}=\mathcal{B}^{-}g^{n+1}+\varphi_{0}(-\tau\mathcal{A}_{\ker})\big{(}u^{n}-\mathcal{B}^{-}g^{n}\big{)}+\tau\varphi_{1}(-\tau\mathcal{A}_{\ker})\big{(}f(t^{n},u^{n})-\mathcal{B}^{-}\dot{g}^{n}\big{)}.$
@@
Consider Assumptions 2.1, 2.2, 2.4 and let $\mathcal{A}$ be symmetric. Then the exponential Euler scheme satisfies the Lipschitz property (2.8).
@@
A direct calculations shows
@@
The practical implementation of the exponential Euler method involves the solution of several saddle point problems. This includes the computation of $x\coloneqq\mathcal{B}^{-}g^{n}\in\mathcal{V}_{c}\subseteq\mathcal{V}$, which can be expressed as the solution of the stationary auxiliary problem
@@
(4.8b) $\mathcal{B}x$ $=$ $\ g^{n}$ in $\qquad\text{in }\mathcal{Q}^{*}.$
The definition of $\varphi_{1}$ implies $\tau\varphi_{1}(-\tau\mathcal{A}_{\ker})\,h=-\big{[}\varphi_{0}({-\tau\mathcal{A}_{\ker}})-\operatorname{id}\big{]}\,\mathcal{A}_{\ker}^{-1}\,h$ for all $h\in\mathcal{H}_{\ker}$. Hence, the exponential Euler scheme can be rewritten as
$u^{n+1}=\mathcal{B}^{-}g^{n+1}+\varphi_{0}(-\tau\mathcal{A}_{\ker})\big{(}u^{n}-\mathcal{B}^{-}g^{n}-w^{n}\big{)}+w^{n},$
@@
To compute the action of $\varphi_{0}(-\tau\mathcal{A}_{\ker})$, one can consider the corresponding PDAE formulation. This leads to
(4.10) $u^{n+1}=\mathcal{B}^{-}g^{n+1}+z(t^{n+1})+w^{n},$
where $z$ is the solution of linear and homogeneous system
@@
on the time interval $[t^{n},t^{n+1}]$ with initial condition $z(t^{n})=u^{n}-\mathcal{B}^{-}g^{n}-w^{n}$.
Adding perturbations to this type of methods is a bit more delicate than for the previously introduced methods. Moreover, there are various possible ways to do this while maintaining the model’s constraint. We discuss two approaches in more detail.
4.3.1. Variant A
@@
where $\sigma$ is again the scaling factor of the noise and the $\tilde{\xi}^{n}$ are i.i.d. standard normal random variables i.e., $\tilde{\xi}^{n}\sim\mathcal{N}(0,1)$. We summarize the necessary steps in Algorithm 1.
Algorithm 1 Probabilistic exponential Euler scheme (Variant A)
@@
3:     compute $\mathcal{B}^{-}g^{n}$, $\mathcal{B}^{-}g^{n+1}$, and $\mathcal{B}^{-}\dot{g}^{n}$ by (4.8)
@@
5:     compute $z$ as the solution of (4.11) on $[t^{n},t^{n+1}]$ with initial data $U^{n}-\mathcal{B}^{-}g^{n}-w^{n}$
@@
In this variant, a single step of the probabilistic exponential Euler scheme involves the solution of four stationary and one transient saddle point problems with only one evaluation of the nonlinear function $f$. Note that all these systems are linear and that the time-dependent system is homogeneous, allowing it to be solved without further regularization.
4.3.2. Variant B
A second possibility is to add the perturbations within the computation of $z$ as right-hand side of (4.11a). For this, we replace the homogeneous system (4.11) by
[…]
(4.13b) $\mathcal{B}z(t)$ $=0$ in $\qquad\text{in }\mathcal{Q}^{*},$
where $\sigma$ and $\tilde{\xi}^{n}$ are defined as in Variant A. The necessary steps for this variant are summarized in Algorithm 2.
Algorithm 2 Probabilistic exponential Euler scheme (Variant B)
2:for $n=0$ to $N-1$ do
3:     compute $\mathcal{B}^{-}g^{n}$, $\mathcal{B}^{-}g^{n+1}$, and $\mathcal{B}^{-}\dot{g}^{n}$ by (4.8)
4:     compute $w^{n}$ by (4.9)
5:     compute $z$ as the solution of (4.13) on $[t^{n},t^{n+1}]$ with initial data $U^{n}-\mathcal{B}^{-}g^{n}-w^{n}$ and parameters $\sigma$, $p=1$, and i.i.d. $\tilde{\xi}^{n}\sim\mathcal{N}(0,1)$
6:     set $U^{n+1}=\mathcal{B}^{-}g^{n+1}+z(t^{n+1})+w^{n}$
7:end for
Remark 4.8.
We would like to emphasize that, in this variant, we use the exponent $p-1/2$ and still obtain the expected order of convergence. This is due to the fact that the perturbation occurs on the right-hand side and is, hence, integrated once, leading to an additional order of $\tau$.
@@ 4.4. Probabilistic exponential integrator of second order
(4.14) $u^{n+1}=u^{n+1}_{\text{Eul}}+\tau\varphi_{2}(-\tau\mathcal{A}_{\ker})\Big{[}f(t^{n+1},u^{n+1}_{\text{Eul}})-\mathcal{B}^{-}\dot{g}^{n+1}-f(t^{n},u^{n})+\mathcal{B}^{-}\dot{g}^{n}\Big{]}$
as introduced in [AZ20]. Therein, $u_{\text{Eul}}^{n+1}$ denotes the result of one step of the (deterministic) exponential Euler scheme (4.10). Considering only the part in the kernel, we get
@@
Consider Assumptions 2.1, 2.2, 2.4 and let $\mathcal{A}$ be symmetric. Then the exponential integrator of second order satisfies the Lipschitz property (2.8).
@@
A direct calculations together with Lemma 4.5 demonstrates
@@
$\qquad\leq\big{(}1+\tau\tilde{L}_{\Psi}\big{)}\,\big{\|}v_{1}-v_{2}\big{\|}_{\mathcal{H}},$
@@
Moreover, $z(t^{n+1})$ is the solution of the homogeneous system (4.11) on the time interval $[t^{n},t^{n+1}]$ with initial value $z(t^{n})=\widehat{w}^{n}$. For more details and the corresponding convergence proof, we refer to [AZ20, Sect. 4].
For the construction of a randomized exponential integrator of second order, we follow Variant A from the previous section but with $p=2$. This then leads to the probabilistic integrator shown in Algorithm 3.
@@
3:     compute one step of the (deterministic) exponential Euler method $U^{n+1}_{\text{Eul}}$
4:     compute $\mathcal{B}^{-}\dot{g}^{n+1}$ by (4.8)
@@ 4.5. Numerical example
For the numerical solution, we have used a second-order finite difference approximation of the Laplacian with $100$ grid points. As time interval, we consider $[0,T]=[0,0.1]$. We vary the value of $p$ in Assumption 2.8, namely $p\in\{1/2,1,3/2\}$ for the first-order methods and $p\in\{1/2,1,3/2,2\}$ for the second-order methods. Then, we compute $1.000$ trajectories of the numerical solution with a fixed noise scale $\sigma=4$ and compute the approximate mean square order of convergence for the implicit Euler method, the midpoint scheme, and the exponential integrators. The results are depicted in Figure 4.1 and show the predicted $\min\{p,q\}$-order convergence.

### v3

@@ 4.2. Probabilistic midpoint scheme
The deterministic midpoint scheme applied to the PDAE (2.1) results in
@@
(4.3b) $\mathcal{B}u^{n+1}\hskip 80.52135pt$ $=\phantom{\tau}g(t^{n+1})\hskip 24.75375pt\text{in }\mathcal{Q}^{*},$
where $f^{n+\frac{1}{2}}=f\left(t^{n}+\frac{\tau}{2},\frac{1}{2}(u^{n}+u^{n+1})\right)$. The result is a second-order scheme, i.e., we get (2.7) with $q=2$. Using again the fact that $u^{n+1}_{c}=\mathcal{B}^{-}g(t^{n+1})$ and restricting equation (4.3a) to test functions in $\mathcal{V}_{\ker}$, we obtain
$\big{(}\operatorname{id}+\tfrac{\tau}{2}\mathcal{A}\big{)}\,u_{\ker}^{n+1}=\big{(}\operatorname{id}-\tfrac{\tau}{2}\mathcal{A}\big{)}\,u^{n}_{\ker}+\tau f\big{(}t^{n}+\tfrac{\tau}{2},\tfrac{1}{2}(u^{n}+u^{n+1})\big{)}-\mathcal{B}^{-}\big{(}g(t^{n+1})-g(t^{n})\big{)}\quad\text{in }\mathcal{V}_{\ker}^{*}$
@@
$\Psi_{\tau}(t,v)=\big{(}\operatorname{id}-\tfrac{\tau}{2}\mathcal{A}\big{)}\,v-\tfrac{\tau}{2}\mathcal{A}\Psi_{\tau}(t,v)+\tau f\big{(}t+\tfrac{\tau}{2},\widetilde{v}(t)\big{)}-\mathcal{B}^{-}\big{(}g(t+\tau)-g(t)\big{)}$
with $\widetilde{v}(t)\coloneqq\frac{1}{2}\big{(}v+\mathcal{B}^{-}g(t)+\Psi_{\tau}(t,v)+\mathcal{B}^{-}g(t+\tau)\big{)}$.
@@
Given $v_{1},v_{2}\in\mathcal{V}_{\ker}$, we define $w_{0}$ and $w_{1}$ as in the proof of Lemma 4.1. It directly follows that $\widetilde{v}_{1}-\widetilde{v}_{2}=\frac{1}{2}(w_{0}+w_{1})$. This, in turn, leads to
@@
$\qquad\qquad+\Big{(}f\big{(}t+\tfrac{\tau}{2},\widetilde{v}_{1}\big{)}-f\big{(}t+\tfrac{\tau}{2},\widetilde{v}_{2}\big{)},w_{0}+w_{1}\Big{)}_{\mathcal{H}}$
@@
$\begin{bmatrix}\operatorname{id}+\frac{\tau}{2}\mathcal{A}&\tau\mathcal{B}^{*}\\ \mathcal{B}&\end{bmatrix}\begin{bmatrix}U^{n+1}\\ \lambda^{n+1}\end{bmatrix}=\begin{bmatrix}(\operatorname{id}-\frac{\tau}{2}\mathcal{A})U^{n}+\tau f^{n+\frac{1}{2}}+\xi^{n}(\tau)\\ g(t^{n+1})\end{bmatrix},$
@@ 4.3. Probabilistic exponential Euler scheme
As a third example, we derive the probabilistic version of the exponential Euler method. The corresponding deterministic scheme for constrained systems has been introduced and examined in [AZ20]. Recall that $\mathcal{A}_{\ker}$ denotes the restriction of $\mathcal{A}$ to the kernel of the constraint operator $\mathcal{B}$. Moreover, we introduce $\mathcal{H}_{\ker}$ as the closure of $\mathcal{V}_{\ker}$ in $\mathcal{H}$. Due to Assumption 2.2, we know that $-\mathcal{A}_{\ker}$ generates an analytic semigroup on $\mathcal{H}_{\ker}$; see [Paz83, Ch. 7, Th. 2.7].
@@
$u_{\ker}(t^{n+1})$ $=u(t^{n+1})-\mathcal{B}^{-}g(t^{n+1})$
(4.4) $=e^{-\tau\mathcal{A}_{\ker}}\big{[}u(t^{n})-\mathcal{B}^{-}g(t^{n})\big{]}+\int_{t^{n}}^{t^{n+1}}e^{-(t^{n+1}-s)\mathcal{A}_{\ker}}\iota_{0}\big{[}f(s,u(s))-\mathcal{B}^{-}\dot{g}(s)\big{]}\,\mathrm{d}s$
with $\iota_{0}\colon\mathcal{H}\equiv\mathcal{H}^{*}\rightarrow\mathcal{H}^{*}_{\ker}\equiv\mathcal{H}_{\ker}$ denoting the restriction of test functions (which will be omitted from now on). Yet another important tool for the construction is the set of the recursively defined $\varphi$-functions,
@@
with $\varphi_{k}(0)=1/k!$ for $z=0$. Since $-\mathcal{A}_{\ker}$ generates an analytic semigroup, it follows from [HO10, Lem. 2.4] that $\varphi_{k}(-\tau\mathcal{A}_{\ker})\colon\mathcal{H}_{\ker}\to\mathcal{H}_{\ker}$, given by
$\varphi_{k}(-\tau\mathcal{A}_{\ker})=-\tau^{-1}\mathcal{A}_{\ker}^{-1}\,\big{(}\varphi_{k-1}(-\tau\mathcal{A}_{\ker})-\varphi_{k-1}(0)\big{)},$
defines a bounded operator.
@@
(4.6) $u^{n+1}=\mathcal{B}^{-}g(t^{n+1})+\varphi_{0}(-\tau\mathcal{A}_{\ker})\big{(}u^{n}-\mathcal{B}^{-}g(t^{n})\big{)}+\tau\varphi_{1}(-\tau\mathcal{A}_{\ker})\big{(}f(t^{n},u^{n})-\mathcal{B}^{-}\dot{g}(t^{n})\big{)}.$
@@
Suppose that Assumptions 2.1, 2.2, and 2.4 hold and let $\mathcal{A}$ be symmetric. Then the exponential Euler scheme satisfies the Lipschitz property (2.8).
@@
Direct calculations show
@@
The practical implementation of the exponential Euler method involves the solution of several saddle point problems. This includes the computation of $x\coloneqq\mathcal{B}^{-}g(t^{n})\in\mathcal{V}_{\text{c}}\subseteq\mathcal{V}$ (see Section 2.3), which can be expressed as the solution of the stationary auxiliary problem
@@
(4.8b) $\mathcal{B}x$ $=$ $\ g(t^{n})$ in $\qquad\text{in }\mathcal{Q}^{*}.$
By [AZ20, Lem. 3.1], this problem yields a unique solution pair $(x,\nu)$ with $\nu$ being a Lagrange multiplier we are not particularly interested in. The definition of $\varphi_{1}$ implies that $\tau\varphi_{1}(-\tau\mathcal{A}_{\ker})\,h=-\big{[}\varphi_{0}({-\tau\mathcal{A}_{\ker}})-\operatorname{id}\big{]}\,\mathcal{A}_{\ker}^{-1}\,h$ for all $h\in\mathcal{H}_{\ker}$. Hence, the exponential Euler scheme can be rewritten as
$u^{n+1}=\mathcal{B}^{-}g(t^{n+1})+\varphi_{0}(-\tau\mathcal{A}_{\ker})\big{(}u^{n}-\mathcal{B}^{-}g(t^{n})-w^{n}\big{)}+w^{n},$
@@
Similarly as before, $\nu^{n}$ acts as Lagrange multiplier corresponding to the constraint. To compute the action of $\varphi_{0}(-\tau\mathcal{A}_{\ker})$, one can consider the corresponding PDAE formulation. This leads to
(4.10) $u^{n+1}=\mathcal{B}^{-}g(t^{n+1})+z(t^{n+1})+w^{n},$
where $z$ is the solution of the linear and homogeneous system
@@
on the time interval $[t^{n},t^{n+1}]$ with initial condition $z(t^{n})=u^{n}-\mathcal{B}^{-}g(t^{n})-w^{n}$. Here, $\mu(t)$ is again a Lagrange multiplier.
Adding perturbations to this type of method is a bit more delicate than for the previously introduced methods. Moreover, there are various possible ways to do this ensuring that the model constraint is satisfied. We discuss one approach in detail.
@@
where $\sigma$ is again the scaling factor of the noise and the $\tilde{\xi}^{n}$ are i.i.d. standard normal random variables i.e., $\tilde{\xi}^{n}\sim\mathcal{N}(0,1)$. The unique solvability of system (4.12) follows again from [AZ20, Lem. 3.1]. We summarize the necessary steps in Algorithm 1.
Algorithm 1 Probabilistic exponential Euler scheme
@@
3:     compute $\mathcal{B}^{-}g(t^{n})$, $\mathcal{B}^{-}g(t^{n+1})$, and $\mathcal{B}^{-}\dot{g}(t^{n})$ by (4.8)
@@
5:     compute $z$ as the solution of (4.11) on $[t^{n},t^{n+1}]$ with initial data $U^{n}-\mathcal{B}^{-}g(t^{n})-w^{n}$
@@
A single step of the probabilistic exponential Euler scheme involves the solution of four stationary and one transient saddle point problems with only one evaluation of the nonlinear function $f$. Note that all these systems are linear and that the time-dependent system is homogeneous, allowing it to be solved without further regularization.
@@ 4.4. Probabilistic exponential integrator of second order
(4.13) $u^{n+1}=u^{n+1}_{\text{Eul}}+\tau\varphi_{2}(-\tau\mathcal{A}_{\ker})\Big{[}f(t^{n+1},u^{n+1}_{\text{Eul}})-\mathcal{B}^{-}\dot{g}(t^{n+1})-f(t^{n},u^{n})+\mathcal{B}^{-}\dot{g}(t^{n})\Big{]}$
as introduced in [AZ20]. Therein, $u_{\text{Eul}}^{n+1}$ denotes the result of one step of the deterministic exponential Euler scheme (4.10). Considering only the part in the kernel, we get
@@
Suppose that Assumptions 2.1, 2.2, and 2.4 hold and let $\mathcal{A}$ be symmetric. Then the exponential integrator of second order satisfies the Lipschitz property (2.8).
@@
Direct calculations together with Lemma 4.6 show
@@
$\qquad\leq\big{(}1+\tau\tilde{L}_{\Psi}\big{)}\,\big{\|}v_{1}-v_{2}\big{\|}_{\mathcal{H}}$
@@
Moreover, $z(t^{n+1})$ is the solution of the homogeneous system (4.11) on the time interval $[t^{n},t^{n+1}]$ with initial value $z(t^{n})=\widehat{w}^{n}$. The Lagrange multipliers $\overline{\nu}^{n}$ and $\widehat{\nu}^{n}$ are once more not of particular interest and simply serve as dummy variables. For more details and the corresponding convergence proof, we refer to [AZ20, Sect. 4].
For the construction of a randomized exponential integrator of second order, we follow the approach from the previous section but with $p=2$. This then leads to the probabilistic integrator shown in Algorithm 2.
@@
3:     compute one step of the deterministic exponential Euler method $U^{n+1}_{\text{Eul}}$ given by (4.10)
4:     compute $\mathcal{B}^{-}\dot{g}(t^{n})$ and $\mathcal{B}^{-}\dot{g}(t^{n+1})$ by (4.8)
@@ 4.5. Numerical example
For the numerical solution, we have used a second-order finite difference approximation of the Laplacian with $100$ grid points. As time interval, we consider $[0,T]=[0,0.1]$. We vary the value of $p$ in Assumption 2.8, namely $p\in\{1/2,1,3/2\}$ for the first-order methods and $p\in\{1/2,1,3/2,2\}$ for the second-order methods. Then, we compute $1000$ trajectories of the numerical solution with a fixed noise scale $\sigma=4$ and compute the approximate mean square order of convergence for the implicit Euler method, the midpoint scheme, and the exponential integrators. The results are depicted in Figure 4.1 and show the predicted $\min\{p,q\}$-order convergence.